OpenStudy (anonymous):

I have to show my work on both sides of the equation 2x -8 =10

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

add 8 to both sides 2x - 8 + 8 = 10 + 8 2x = 18 divide both sides by 2 x = 9

OpenStudy (anonymous):